OpenStudy (anonymous):

construct and complete a truth table with the following headings: 1) p 2) q 3) p → q 4) p ∧ (p → q) 5) [p ∧ (p → q)] → q

OpenStudy (anonymous):

we can do this

OpenStudy (anonymous):

if you are here that is, it is not that hard

OpenStudy (anonymous):

ready?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I have no idea how to do this lol @satellite73

OpenStudy (anonymous):

ok first we start here \[\begin{array}{c|c} p & q \\ \hline & \\ & \\ & \\ & \\ \end{array}\]

OpenStudy (anonymous):

then we put all combinations of t and f underneath usually like this \[\begin{array}{|c|c} p & q \\ \hline T & T \\ T & F \\ F & T \\ F & F \\ \hline \end{array}\]

OpenStudy (anonymous):

that that each possible pair is there have you seen something like this before?

OpenStudy (anonymous):

next step is \[\begin{array}{|c|c|c} p & q & p\to q \\ \hline T & T & \\ T & F & \\ F & T & \\ F & F & \\ \hline \end{array}\]

OpenStudy (anonymous):

\(p\to q\) is always true, unless \(p\) is true and \(q\) is false it will look like this \[\begin{array}{|c|c|c} p & q & p\to q \\ \hline T & T &T \\ T & F & F\\ F & T & T\\ F & F & T\\ \hline \end{array}\]

OpenStudy (anonymous):

how we doing so far? are you totally lost, or is it okay ?

OpenStudy (anonymous):

oh okay im starting to get it

OpenStudy (anonymous):

the first two columns are there to get all possible combinations of \(T\) and\(F\) for \(p,q\)

OpenStudy (anonymous):

the third column \(p\to q\) as i said is always true unless \(p\) is \(T\) and \(q\) is \(F\)

OpenStudy (anonymous):

as your read across two more to go

OpenStudy (anonymous):

This is what the project says: On a sheet of paper, construct and complete a truth table with the following headings: p q p → q p ∧ (p → q) [p ∧ (p → q)] → q 2. Construct and complete a truth table in order to determine if the following statements are logically equivalent: p → q and ~q → ~p 3. Explain in the box below if the statement [p ∧ (p → q)] → q is a tautology or not. 4. Explain in the box below if the statements p → q and ~q → ~p are logically equivalent or not. 5. Turn your paper with your truth tables in to your teacher.

OpenStudy (anonymous):

\[\begin{array}{|c|c|c|c|} p & q & p\to q &p\land (p\to q) \\ \hline T & T &T & \\ T & F & F&\\ F & T & T&\\ F & F & T&\\ \hline \end{array}\]

OpenStudy (anonymous):

we will get there if you have the patience for it we are not there yet

OpenStudy (anonymous):

we have to fill in the column \(p\land (p\to q)\)

OpenStudy (anonymous):

for this column you only look at the column under \(p\) and the column under \(p\to q\) the "and" statement is only true if there is a T in both rows

OpenStudy (anonymous):

\[\begin{array}{|c|c|c|c|} p & q & p\to q &p\land (p\to q) \\ \hline T & T &T &T \\ T & F & F&F\\ F & T & T&F\\ F & F & T&F\\ \hline \end{array}\]

OpenStudy (anonymous):

if i have totally lost you let me know or if you have a specific question please ask we have one more column to fill

OpenStudy (anonymous):

What does the ^ mean

OpenStudy (anonymous):

\(p\land q\) means "\(p\) AND \(q\)" it is only True if both \(p\) and \(q\) are True

OpenStudy (anonymous):

like saying "it is raining and cold" is only true if it is both raining and cold

OpenStudy (anonymous):

okay but i dont get what the last column means all together. p ∧ (p → q)

OpenStudy (anonymous):

in english it is kind of funny it says "p and p implies q"

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

if we change p to "it is sunny" and q to "i go to the beach" it says "it is sunny and if it is sunny i will go to the beach"

OpenStudy (anonymous):

oh okay!

OpenStudy (anonymous):

last column ready? it is going to take me a minute to format it

OpenStudy (anonymous):

okay ready

OpenStudy (anonymous):

\[\begin{array}{|c|c|c|c|c} p & q & p\to q &p\land (p\to q) &[p\land (p\to q)]\to q \\ \hline T & T &T &T& \\ T & F & F&F&\\ F & T & T&F&\\ F & F & T&F&\\ \hline \end{array}\]

OpenStudy (anonymous):

the reason they have you do one at a time is so we can understand the last one the last one is the \([p\land (p\to q)]\to q\) which is always true unless \(p\land (p\to q)\) is true and \(q\) is false so we look at the column under \(p\land (p\to q)\) and \(q\) and if we see \(T, F\) we put \(F\) otherwise we put \(T\)

OpenStudy (anonymous):

but if you look carefully, you see that it is never the case if \(p\land (p\to q)]\) is \(T\) then \(q\) is also \(T\) in that row \[\begin{array}{|c|c|c|c|c} p & q & p\to q &p\land (p\to q) &[p\land (p\to q)]\to q \\ \hline T &\color{red} T &T &\color{red}T& \\ T & F & F&F&\\ F & T & T&F&\\ F & F & T&F&\\ \hline \end{array}\]

OpenStudy (anonymous):

therefore the last column gets all T\[\begin{array}{|c|c|c|c|c} p & q & p\to q &p\land (p\to q) &[p\land (p\to q)]\to q \\ \hline T & T &T &T& T\\ T & F & F&F&T\\ F & T & T&F&T\\ F & F & T&F&T\\ \hline \end{array}\]

OpenStudy (anonymous):

now we can answer question 3 as well since the last column has all \(T\) that means it is always true, and is a "tautology"

OpenStudy (anonymous):

back to the previous example, the last column says "it is sunny, and if it is sunny i go to the beach, then i go to the beach" pretty obviously true right?

OpenStudy (anonymous):

yeah so they are equivalent?

OpenStudy (anonymous):

hold the phone you need two different statements to ask "are they equivalent" we were answering this question 3. Explain in the box below if the statement [p ∧ (p → q)] → q is a tautology or not.

OpenStudy (anonymous):

oh i thought were on 4 lol

OpenStudy (anonymous):

the answer to that one is "yes, it is a tautology"

OpenStudy (anonymous):

#4 requires doing #2, which is a lot easier than the last one

OpenStudy (anonymous):

okay how do i figure it out

OpenStudy (anonymous):

ok here we go

OpenStudy (anonymous):

\[\begin{array}{|c|c|c|c|c|c} p & q & \lnot{}p & \lnot{}q & p\to q & \lnot{}q\to\lnot{}p \\ \hline T & T & & & & \\ T & F & & & & \\ F & T & & & & \\ F & F & & & & \\ \hline \end{array}\]

OpenStudy (anonymous):

as before, first two columns give all combinations of T and F now to get \(\lnot p\) and \(\lnot q\) change all T to F and F to T

OpenStudy (anonymous):

do you know what \(\lnot p\) means?

OpenStudy (anonymous):

isnt it just the opposite of p?

OpenStudy (anonymous):

\[\begin{array}{|c|c|c|c|c|c} p & q & \lnot{}p & \lnot{}q & p\to q & \lnot{}q\to\lnot{}p \\ \hline T & T & F &F \\ T & F &F & T& & \\ F & T & T& F & & \\ F & F & T& T& \\ \hline \end{array}\] yes it means "not p"

OpenStudy (anonymous):

you see how i got the column \(\lnot p\)? i looked at the column for \(p\) and if i see T i put F and if i see F i put T same for \(\lnot q\)

OpenStudy (anonymous):

now we already did \(p\to q\) for the last question it is always T unless \(p\) is T and \(q\) is F, in the second row

OpenStudy (anonymous):

\[\begin{array}{|c|c|c|c|c|c} p & q & \lnot{}p & \lnot{}q & p\to q & \lnot{}q\to\lnot{}p \\ \hline T & T & F &F &T \\ T & F &F & T&F & \\ F & T & T& F & T& \\ F & F & T& T& T \\ \hline \end{array}\]

OpenStudy (anonymous):

and finally we look at \(\lnot q\to \lnot p\) which is always true unless \(\lnot q\) it T and \(\lnot p\) is F

OpenStudy (anonymous):

that only occurs here \[\begin{array}{|c|c|c|c|c|c} p & q & \lnot{}p & \lnot{}q & p\to q & \lnot{}q\to\lnot{}p \\ \hline T & T & F &F &T \\ T & F &\color{red}F & \color{red}T&F & \\ F & T & T& F & T& \\ F & F & T& T& T \\ \hline \end{array}\]

OpenStudy (anonymous):

fill it in like this \[\begin{array}{|c|c|c|c|c|c} p & q & \lnot{}p & \lnot{}q & p\to q & \lnot{}q\to\lnot{}p \\ \hline T & T & F &F &T&T \\ T & F &\color{red}F & \color{red}T&F &F \\ F & T & T& F & T& T \\ F & F & T& T& T &T\\ \hline \end{array}\]

OpenStudy (anonymous):

now we look at the last two columns and see that they are identical guess what that means?

OpenStudy (anonymous):

they are equivalent!

OpenStudy (anonymous):

exactly back to our english example if it is sunny, i will go to the beach if i did not go to the beach, then it is not sunny they say the same thing

OpenStudy (anonymous):

Oh okay could you help me set up the chart?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

set up what chart?

OpenStudy (anonymous):

Oh nvm you already did i think for number 2 it says to set up the truth table

OpenStudy (anonymous):

lordamercy what do you think we were doing for the past hour?!

OpenStudy (anonymous):

lol but which chart is the one that proves they are equivalent

OpenStudy (anonymous):

the entire table is the truth table for both the fact that the last two columns are identical is what tells you the two statements are equivalent

OpenStudy (anonymous):

\[\begin{array}{|c|c|c|c|c|c} p & q & \lnot{}p & \lnot{}q & p\to q & \lnot{}q\to\lnot{}p \\ \hline T & T & F &F &T&T \\ T & F &F & T&F &F \\ F & T & T& F & T& T \\ F & F & T& T& T &T\\ \hline \end{array}\]this is the truth table for both \(p\to q\) and \(\lnot q\to \lnot p\) under both, you see the same thing